# Xuanyu Zhang2,Qing Yang2, Dongliang Xu2

### v2

@@ 4.2 Main Results
Tokens/s Speed
AR 54.98 1.00
Random 95.07 1.73
Zero 102.68 1.87
Fixed 117.43 2.12
Shuffle 118.78 2.16
TR 119.56 2.17
@@ 5.2 Ablation Study
Section 3.1 discuss duplicate tokens in the merged sequence during matrix updates. We compare three updating strategies: using candidate tokens from the first occurrence, from the last occurrence, and the current method (merging via parallel CUDA operations). Figure 3(c) indicates that using the last occurrence yields the highest MAT, which may benefit from more contextual information. However, the differences among different strategies in MAT are minimal. In terms of Tokens/s, the current approach significantly outperforms the other two, as it avoids the additional processing required to manage token positions, thereby reducing delays. Speculative decoding is highly sensitive to latency, any extra operation must provide substantial benefits to outweigh its time cost.
@@ 6 Related Work
Efficient inference is crucial for real-time applications and low-resource scenarios. Many strategies have been developed to reduce latency Zhou et al. (2024b). Among these, speculative decoding Chen et al. (2023); Leviathan et al. (2023); Miao et al. (2024); Xia et al. (2023) is a losses technique that predicts multiple possible continuations simultaneously. It reduces the number of decoding steps needed without compromising accuracy. Some speculative decoding methods rely on additional draft models to guess draft tokens. These typically involve using smaller models from the same series Zhao et al. (2024); Spector and Re ; Sun et al. (2023); Liu et al. (2024b); Yuan et al. (2024); Gong et al. (2024) or training new models with a shared vocabulary Leviathan et al. (2023); Chen et al. (2023); Zhou et al. (2024a); Li et al. (2024). It is worth noting that Zhao et al. (2024) also uses rejected tokens but does not include candidate tokens. Additionally, Kou et al. (2024); Wang et al. (2024b) propose training the original LLMs to enable non-aggressive decoding. While effective, these approaches require managing or training multiple models, which can be non-trivial and resource-intensive. Other methods focus on parameter-efficient structures. These approaches minimize the need for complete retraining but still require model-specific training and adaptation, limiting their scalability and general applicability Lin et al. (2024); Liu et al. (2024a).
@@
Token Recycling is a train-free, retrieval-based method. It requires no additional generation, covers a broader range of possible continuations, and demands minimal storage with low retrieval costs. The update process ensures the retrieval space remains adaptable.
@@ A.2 Draft Tree Algorithm and Structure
In Token Recycling, we also use a heuristically constructed token tree to perform the verification. As shown in Figure 5, we construct a static and unbalanced tree inspired by  Cai et al. (2024). The number $k$ on a node indicates that it is the $k$-th candidate token for its parent node. The construction process is below. We begin with a fully balanced 10-branch tree and use an independent validation set to identify the top $k$ nodes that most frequently yield correct tokens. These top $k$ nodes and their children are retained to form a new tree, and the process is repeated to identify the next set of top $k$ nodes. This iterative process continues until performance no longer shows significant improvement. The final tree is determined, and the $k$ is set to consider the maximum number of children across all nodes and the memory requirement. While empirical, this iterative approach has proven to be effective. Further details on tuning the $n$ are provided in Section 5.1. Overall, the tree we construct contains 80 nodes (including the root node) in 6 layers. This means that each forward requires an additional draft input of 79 tokens with a maximum acceptance length of 6.
@@
A.3 Time Allocation
For speculative decoding to be effective, it is essential to maintain a high hit rate while minimizing the time spent on additional operations. We divide each decoding step into several components: preprocessing , retrieving draft tokens, model forward pass, verifying draft sequences, and updating the matrix, input tokens, and key-value cache. The average time spent on each component is shown in Figure 4. The results indicate that the majority of the time is consumed by the model forward pass. The verification process also takes a significant amount of time due to the need to extract and verify all feasible paths. Retrieving draft tokens and updating operations take roughly the same amount of time.
A.4 Compare With Eagle
TR is Self-Draft and Train-Free, and fair comparisons with other speculative decoding methods that do not require training (e.g., Lookahead, PLD, REST) have already been conducted in section 4.2. However, due to the effectiveness of our approach, our performance significantly exceeds that of these baselines. As a result, we became curious about comparing our method with training-dependent approaches, such as representively, Medusa. The comparison results show that our method even outperforms it a lot. Approaches like Eagle Li et al. (2024), which rely on additional architectures to generate drafts, may achieve higher performance, but they require collecting datasets for training the small model, and during deployment, they also need extra memory. In contrast, TR only requires <2MB of storage, is train-free, and can be directly applied to all LLMs, achieving nearly a 2x speedup across various model sizes. In Table 6, we compare the performance of Token Recycling and Eagle on the vicuna-7b-v1.3 model using SpecBench. In comparison, our method uses only 0.39% of the memory required by Eagle, while achieving 88.65% of its speedup performance.
Memory Cost Speed Ratio
Eagle 500MB 2.29
TR <2MB 2.03
Table 6: Compare Token Recycling with Eagle.
A.5 Reuse Mechanism Analysis
@@
Rejected tokens may not satisfy the long-range dependencies of the previous, but this does not mean they do not meet the long-range dependencies of the current. These tokens may be accepted in the future. As shown in Table 4, the performance gain from the candidate tokens of rejected tokens is significantly greater than the gain from using only the accepted tokens in the SpecBench experiments. We also include a case study in A.6.
A.6 Case Study

### v3

@@ 5.2 Ablation Study
Tokens/s Speed
AR 54.98 1.00
Random 95.07 1.73
Zero 102.68 1.87
Fixed 117.43 2.12
Shuffle 118.78 2.16
TR 119.56 2.17
@@
Section 3.1 discusses duplicate tokens in the merged sequence during matrix updates. We compare three updating strategies: using candidate tokens from the first occurrence, from the last occurrence, and the current method (merging via parallel CUDA operations). A detailed explanation of three strategies is in Appendix A.5 Figure 3(c) indicates that using the last occurrence yields the highest MAT, which may benefit from more contextual information. However, the differences among different strategies in MAT are minimal. In terms of Tokens/s, the current approach significantly outperforms the other two, as it avoids the additional processing required to manage token positions, thereby reducing delays. Speculative decoding is highly sensitive to latency, any extra operation must provide substantial benefits to outweigh its time cost.
@@
Figure 4: MAT and Speedup ratio under different temperatures during generation.
Temperature Sampling
To enhance diversity, a temperature greater than 0 is often employed during LLM generation. We conducted experiments on SpecBench at different temperature settings to investigate the impact of this randomness on acceleration. Figure 4 shows that at a temperature of 0.3, the performance remains unaffected when compared to greedy decoding. However, as the temperature increases to 0.5 and 1, a slight performance degradation is observed. This is likely due to we only store the top-k candidate tokens. At higher sampling temperatures, the probability of selecting tokens outside the top-k increases. Despite this, the speedup ratio of TR remains consistently around 2.0. The sustained acceleration performance under various sampling settings demonstrates the robustness and broad applicability of TR.
Method Memory (MB) Tokens/s Speed
Eagle1 500 106.94 2.08
Eagle2 500 116.95 2.28
TR 1.95 107.52 2.09
Table 5: The comparison of Eagle1/2 with TR about memory costs and speedup ratio.
5.3 Compare with Eagle
Table 1 demonstrates that TR significantly outperforms the train-need Medusa. We were curious to compare TR with the state-of-the-art speculative decoding methods - Eagle1 Li et al. (2024b) and Eagle2 Li et al. (2024a). Eagle1 is a train-dependent method that collects training data and trains a separate draft model using the hidden states and token embeddings from a large LLM. Eagle2 improves upon Eagle1 by incorporating dynamic trees into the draft tree construction process. In Table 5, TR is compared with Eagle1/2 on SpecBench. It is a surprise that, despite being completely training-free and self-drafting, TR outperforms Eagle1. Moreover, TR achieves 91.23% of the acceleration performance of Eagle2 while only requiring 0.39% of the memory used by Eagle2. It is worth noting that TR still employs a static tree, rather than the dynamic tree used in Eagle2. This highlights the remarkable effectiveness and efficiency of TR and shows potential for further improvement when combine TR with dynamic trees.
5.4 Time Allocation
Effective speculative decoding requires not only a high hit rate but also minimal time on additional operations. Each decoding step can be divided into several components: preprocessing , retrieving draft tokens, model forward, verifying draft sequences, and updating the matrix, input tokens, and key-value cache. Figure 5 shows that the majority of the time is consumed by the model forward. The second most is verification which involves extracting and validating all feasible paths. In contrast, TR introduces only negligible latency in its dedicated preprocessing, retrieval, and update steps. This highlights the efficiency of TR’s design.
@@ 6 Related Work
Efficient inference is crucial for real-time applications and low-resource scenarios. Many strategies have been developed to reduce latency Zhou et al. (2024b). Among these, speculative decoding Chen et al. (2023); Leviathan et al. (2023); Miao et al. (2024); Xia et al. (2023) is a losses technique that predicts multiple possible continuations simultaneously. It reduces the number of decoding steps needed without compromising accuracy. Some speculative decoding methods rely on additional draft models to guess draft tokens. These typically involve using smaller models from the same series Zhao et al. (2024); Spector and Re ; Sun et al. (2023); Liu et al. (2024b); Yuan et al. (2024); Gong et al. (2024) or training new models with a shared vocabulary Leviathan et al. (2023); Chen et al. (2023); Zhou et al. (2024a); Li et al. (2024b). It is worth noting that Zhao et al. (2024) also uses rejected tokens but does not include candidate tokens. Additionally, Kou et al. (2024); Wang et al. (2024b) propose training the original LLMs to enable non-aggressive decoding. While effective, these approaches require managing or training multiple models, which can be non-trivial and resource-intensive. Other methods focus on parameter-efficient structures. These approaches minimize the need for complete retraining but still require model-specific training and adaptation, limiting their scalability and general applicability Lin et al. (2024); Liu et al. (2024a).
@@
TR is a train-free, retrieval-based method. It requires no additional generation, covers a broader range of possible continuations, and demands minimal storage with low retrieval costs. The update process ensures an adaptable retrieval space.
@@ 7 Conclusion
Acknowledge
We gratefully acknowledge the support of the National Natural Science Foundation of China (NSFC) via grant 62236004, 62206078 and 62476073.
@@ A.2 Draft Tree Algorithm and Structure
In Token Recycling, we also use a heuristically constructed token tree to perform the verification. As shown in Figure 6, we construct a static and unbalanced tree inspired by  Cai et al. (2024). The number $k$ on a node indicates that it is the $k$-th candidate token for its parent node. The construction process is below. We begin with a fully balanced 10-branch tree and use an independent validation set to identify the top $k$ nodes that most frequently yield correct tokens. These top $k$ nodes and their children are retained to form a new tree, and the process is repeated to identify the next set of top $k$ nodes. This iterative process continues until performance no longer shows significant improvement. The final tree is determined, and the $k$ is set to consider the maximum number of children across all nodes and the memory requirement. While empirical, this iterative approach has proven to be effective. Further details on tuning the $n$ are provided in Section 5.1. Overall, the tree we construct contains 81 nodes (including the root node) in 6 layers. This means that each forward requires an additional draft input of 79 tokens with a maximum acceptance length of 6.
@@
A.3 Reuse Mechanism Analysis
@@
Rejected tokens may not satisfy the long-range dependencies of the previous, but this does not mean they do not meet the long-range dependencies of the current. These tokens may be accepted in the future. As shown in Table 4, the performance gain from the candidate tokens of rejected tokens is significantly greater than the gain from using only the accepted tokens in the SpecBench experiments. We also include a case study in A.4.
A.4 Case Study
@@
A.5 Explanation of Three Update Strategies
Suppose our merged sequence is:
$S=[\underline{a},b,c,\underline{a},d]$
, which token $a$ appears twice. Currently, corresponding output tokens are stored in matrix $O$:
$O[0]$ $=[a_{1},a_{2},a_{3},a_{4}]$
$O[1]$ $=[b_{1},b_{2},b_{3},b_{4}]$
$O[2]$ $=[c_{1},c_{2},c_{3},c_{4}]$
$O[3]$ $=[a_{5},a_{6},a_{7},a_{8}]\text{ (second occurrence of }a\text{)}$
$O[4]$ $=[d_{1},d_{2},d_{3},d_{4}]$
The three strategies for updating the adjacency matrix $M$ are:
  • Cur (uncontrolled): Update directly without position control, potentially mixing outputs from repeated tokens. For example: $M[S]=O,M[a]=[a_{1},a_{2},a_{3},\underline{a_{8}}]$ (the last $\underline{a_{8}}$ is from the second occurrence).
  • First: Record positions of each token’s first occurrence: $\text{pos}=[0(a),1(b),2(c),4(d)]$. Resulting in: $M[S[pos]]=O[pos],M[a]=[a_{1},a_{2},a_{3},a_{4}]$
  • Last: Record positions of each token’s last occurrence: $\text{pos}=[1(b),2(c),3(a),4(d)]$. Resulting in: $M[S[pos]]=O[pos],M[a]=[a_{5},a_{6},a_{7},a_{8}]$
These different update strategies have a slight impact on Mean Accepted Tokens. However, extracting these positions introduces additional latency, ultimately reducing overall acceleration effectiveness compared to the simpler uncontrolled approach.